OpenStudy (anonymous):

[7.04] Choose the correct simplification of (3a3b2c)0. 0 3a3b2c 1 3

OpenStudy (anonymous):

Anything times 0 is 0. (3a x 3b x3c) x 0 = 0

OpenStudy (anonymous):

anything to the power of zero is one

OpenStudy (anonymous):

so is it o or 1

OpenStudy (anonymous):

\[0 \times 0 = 0\] \[19 \times 0 = 0 \]

OpenStudy (anonymous):

If it's a power it's a 1 like satellite said but if it's just a 0 the answer is 0.

OpenStudy (anonymous):

\[ (3a3b2c)^0=1\]

OpenStudy (anonymous):

\[198^0 = 1\]

OpenStudy (anonymous):

i am going to guess the problem is \[ \huge(3a^3b^2c)^0 \]

OpenStudy (anonymous):

^ yeah me too

OpenStudy (anonymous):

you guys are telling me its 0

OpenStudy (anonymous):

in which case, since anything raised to the power of zero is one, the answer is one

OpenStudy (anonymous):

users do not usually take the time and fix the equations. i.e the exponents and such.

OpenStudy (anonymous):

@Felicia123 one one it is one

OpenStudy (anonymous):

It's one. He's right.

OpenStudy (anonymous):

\[1\]

OpenStudy (anonymous):

Oh ok s its one not 0

OpenStudy (anonymous):

lol. omg does \[1 ~ look ~ like ~ 0\]

OpenStudy (anonymous):

Yeah it's one. I thought it was times 0 not to the power of 0. So it is one.

OpenStudy (anonymous):

\[The ~ Answer ~ is ~ simply ~ 1 ~ :)\]

OpenStudy (anonymous):

omg lol thank you

OpenStudy (anonymous):

lol no prob